OpenStudy (anonymous):

the length of rectangle is twice the width;the perimeter is 42 inches, find the width? someone please help i will really appreciate it and if you are able to explain how to do it :)

OpenStudy (anonymous):

Write the formula for perimeter of a rectangle. Then, since: "the length of rectangle is twice the width" you can replace the length with twice the width. So instead of L, write 2w. Then, when you simplify, you'll have an equation with just w as an unknown. Solve for w and you're done. Try it and if you have problems say where and you'll get help.

OpenStudy (anonymous):

the formula is P=2(L) + 2(W) i dont understand what is it that i have to replace

OpenStudy (anonymous):

replace L with 2*W because it says: "the length is twice the width" Now your equation is: P=2(2W)+2(W) Also, you know P=42 42=2*2*W+2*W simplify and solve for W.

OpenStudy (anonymous):

p=2(L) + 2(W) 42=2(2W) + 2(w) 42=4w + 2w 42= 6w 42/6=6/6 7=W is this correct and do i have any other step like plugging it in